OpenStudy (anonymous):

int between pi and -pi (sin5xcox3x)

OpenStudy (anonymous):

\[\int\limits_{-\Pi}^{\Pi}\sin 5x \cos 3x\]

OpenStudy (anonymous):

use trig identities for sin(u+v) and cos(u+v) and expand out to sin(x) and cos(x)

OpenStudy (anonymous):

I understand the trig ident thus giving answer of -1/16 cos 8x -1/4 cos2x... but if its between pi and -pi isn't zero?

OpenStudy (anonymous):

cos(x)=cos(-x)

OpenStudy (anonymous):

cos pi and cos -pi both equal -1 ?

OpenStudy (anonymous):

so you have (1/16+1/4)*2

OpenStudy (anonymous):

I have tried two different ways and get zero both ways. yes, cos(PI)=cos(-PI)=-1. That is why on the interval (-PI,PI), it yields a zero.